OpenStudy (anonymous):

m=-1;(-3,6) write a point slpope equation for the line with the given slope that centers the given point y-6=-1(x-3) this my answer

OpenStudy (sriram):

y-y'=m(x-x') y-6=-1(x-(-3)) =-1(x+3)

OpenStudy (anonymous):

ok so i need a + sign then and not a - sign

OpenStudy (sriram):

yep